# A Survey on Mixture of Experts

### v1

@@ 7. Challenges & Opportunities
Mixture of Experts (MoE) models present a compelling approach for significantly increasing model capacity at a constant computational cost. Despite their promise, several intrinsic challenges remain. In this section, we identify critical challenges and promising directions for future investigation as follows:
@@
Training Stability and Load Balancing. MoE models that utilize sparse gating have become a popular means to expand model capacity without proportionally increasing computational demands. However, the discrete nature of assigning a fixed number of experts to tokens leads to significant challenges in maintaining balanced workloads of experts and training stability across varying inputs (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022b; Zhou et al., 2022a; Antoniak et al., 2023). Load imbalances, where certain experts become over-utilized while others are underutilized can hinder expert specialization and further degrade model performance. Although current efforts (Lepikhin et al., 2020; Fedus et al., 2022b; Du et al., 2022; Jiang et al., 2024a; Lieber et al., 2024; Dai et al., 2024; Tianwen Wei, 2024) have attempted to address this challenge by incorporating auxiliary loss functions to encourage even token distribution across experts, these solutions can still lead to training instability (Zoph et al., 2022) and often neglect the relative importance of different tokens (Zhou et al., 2022a). Therefore, future studies should focus on more effective regularization techniques (Zoph et al., 2022) or innovative gating algorithms (Zhou et al., 2022a; Antoniak et al., 2023; Muqeeth et al., 2023; Zhong et al., 2024) that encourage equitable load distribution among experts and enhance model training stability.
@@
Scalability and Communication Overhead. As the escalating sizes of LLMs with MoE necessitate more expansive distributed systems, the imperative for efficient communication during model training becomes increasingly critical, as elaborated in Section 5.2. The trade-off between model complexity, indicated by the number of parameters, and the communication overhead represents a significant bottleneck in distributed training processes (Lepikhin et al., 2020). To address these challenges, it is essential to develop and implement effective strategies that enhance the efficiency of information transfer from system aspect or streamline information exchange without compromising model performance from algorithm aspect. Innovations such as DeepSpeed (Rajbhandari et al., 2022), FasterMoE (He et al., 2022), and ScMoE (Cai et al., 2024) are at the forefront of minimizing communication overhead. For example, the shared expert approach (Rajbhandari et al., 2022; Dai et al., 2024; Team, 2024b, c), advancing MoE with parameter-sharing frameworks, holds promise for reducing the volume of data transmitted between distributed systems while concurrently enhancing model performance in natural language processing tasks. Such innovations are pivotal in facilitating more scalable and efficient distributed training architectures for MoE models.
@@
Expert Specialization and Collaboration. Expert specialization refers to the concept where each expert develops non-overlapping and focused knowledge. Encouraging experts to concentrate their skills on distinct sub-tasks or domains has been shown to enhance the performance and generalization of the MoE model. The prevailing strategy involves designating a select number of experts as shared ones, with the goal of capturing commonalities in knowledge and reducing redundancy among those experts that are routed dynamically (Rajbhandari et al., 2022; Xue et al., 2024; Team, 2024b; Dai et al., 2024). However, fostering effective collaboration among these specialized experts is an ongoing challenge. Relying solely on a sparsely computed weighted sum of outputs from the top- experts can overlook the intricate internal relationships that exist across the entire experts. Consequently, exploring new mechanisms for enhancing both the specialization and collaboration among experts is crucial for the development of more integrated and powerful MoE models.
@@
Sparse Activation and Computational Efficiency. One of the primary benefits of MoE models lies in their capacity for sparse activations, which theoretically enhances computational efficiency. Nevertheless, implementing this efficiency in practice poses substantial challenges. This is attributed to the non-uniformity of sparse operations within hardware accelerators (Dao et al., 2022; Kwon et al., 2023). Furthermore, optimizing the balance between activating a select top- subset of experts from an entire pool of experts entails intricate coordination. This optimization is crucial for ensuring that each expert develops a specialized niche (Dai et al., 2024). Thus, there is a pressing need for further research into hardware optimization techniques that more adeptly accommodate sparse computations. Such advancements would not only preserve the model’s capacity but could also significantly enhance the performance and efficiency of MoE models.
@@
Generalization and Robustness. MoE models have demonstrated increased computational efficiency during pre-training phases. However, there is a notable propensity for sparse MoE architectures to overfit to specific tasks or datasets, which undermines their ability to generalize effectively (Fedus et al., 2022b; Zoph et al., 2022; Shen et al., 2023a; Dou et al., 2023a). To enhance the generalization and robustness of MoE models when encountering unseen data and diverse input variations, various strategies have been explored. These include regularization techniques such as dropout (Fedus et al., 2022b) and token dropping (Zoph et al., 2022), as well as multi-task instruction tuning (Shen et al., 2023a; Dou et al., 2023a). Looking ahead, there is potential for further advancements in this challenge. Future endeavors could explore innovative regularization methods, refined multi-task learning frameworks, or the incorporation of meta-learning concepts that bolster the MoE models’ robustness and extend their generalization capabilities across an even broader spectrum of downstream tasks.
@@
Interpretability and Transparency. The inherent complexity of MoE models, coupled with their dynamic gating of inputs to specialized experts, poses significant challenges to interpretability. This becomes particularly problematic in contexts where comprehending the rationale behind the model’s decisions is essential. Enhancing the interpretability of MoE models is therefore critical, not only to facilitate a clearer understanding of their decision-making processes but also to address underlying challenges such as load balancing (Lepikhin et al., 2020; Fedus et al., 2022b; Du et al., 2022) and the mitigation of knowledge redundancy (Team, 2024b; Dai et al., 2024). In light of these considerations, there is a pressing need for future studies focused on the development of methods and tools that can effectively visualize and explain the behavior of individual experts within MoE models, as well as the nature of their interactions. Such advancements would significantly improve our grasp of MoE models and bolster their ongoing development, ensuring their gating decisions are transparent and trustworthy.
@@
Optimal Expert Architecture. The design of MoE architectures, encompassing the selection of network types and the quantity of experts, significantly influences the efficacy of multi-task learning across various domains. A plethora of network architectures has been adopted as experts, including LSTM (Shazeer et al., 2017), CNN (Chowdhury et al., 2023; Zhang et al., 2023), FFNs (MLPs) (Lepikhin et al., 2020; Fedus et al., 2022b; Zoph et al., 2022; Pan et al., 2024), Attention (Zhang et al., 2022b; Shen et al., 2023c), and LoRA (Dou et al., 2023b; Li et al., 2024b; Luo et al., 2024). Among these, FFNs as experts remain the most prevalent. Despite their considerable achievements, the exploration of various hybrids of network types within experts (as the distinct features processing capabilities of different network architectures), as well as the development of innovative expert architectures, remains nascent areas of research. Furthermore, the strategic allocation of a varying number of experts across different layers of the model presents an area ripe for investigation. This is due to two primary considerations: 1) different layers of the model capture semantic information at varying levels of granularity; 2) an excessive number of experts can complicate the training process and augment computational costs, while an insufficient number of experts might lead to knowledge redundancy and diminish the specialization capabilities of the experts. To navigate these challenges, the development of automated architecture search methods specifically designed for MoE models is imperative (Zhou et al., 2023). Such approaches could systematically identify optimal configurations, balancing the trade-offs between computational efficiency and the specialization of experts.
@@
Integration with Existing Frameworks. Ensuring seamless integration of MoE models into existing large language models (LLMs) is crucial for their broad adoption. It is particularly vital to enable adaptation of LLMs to MoE architecture without necessitating training from scratch, as it can significantly reduce resource consumption. Recent studies (Dou et al., 2023b; Chen et al., 2024; Li et al., 2024b; Luo et al., 2024; Zadouri et al., 2023; Gao et al., 2024; Wu et al., 2024a) have demonstrated the efficacy of combining Parameter-efficient Fine-tuning (PEFT) techniques with MoE frameworks, offering a promising method for incorporating MoE into established LLMs. However, these methods may compromise model performance or complicate the existing parallel strategies of pretraining and inference efforts (Han et al., 2024). Advancing the development of modular and plug-and-play MoE components is essential. Additionally, optimizing these components for training and deployment across diverse computing environments and hardware platforms will expand their applicability. Such advancements are expected to enhance the versatility and efficiency of MoE models, making them more accessible for a wide range of applications and platforms.

### v2

@@ 7 Challenges & Opportunities
Mixture of Experts (MoE) models present a compelling approach for significantly increasing model capacity at a constant computational cost. Despite their promise, several intrinsic challenges remain, necessitating further collaborative design and engineering across algorithm, system, and application aspects. In this section, we identify critical challenges and promising directions for future investigation as follows:
@@
Training Stability and Load Balancing. MoE models that utilize sparse gating have become a popular means to expand model capacity without proportionally increasing computational demands. However, the discrete nature of assigning a fixed number of experts to tokens leads to significant challenges in maintaining balanced workloads of experts and training stability across varying inputs [27, 28, 36, 94, 56]. Load imbalances, where certain experts become over-utilized while others are underutilized can hinder expert specialization and further degrade model performance. Although current efforts [28, 36, 35, 29, 68, 69, 67] have attempted to address this challenge by incorporating auxiliary loss functions to encourage even token distribution across experts, these solutions can still lead to training instability [37] and often neglect the relative importance of different tokens [94]. Therefore, future studies should focus on more effective regularization techniques [37] or innovative gating algorithms [94, 56, 40, 41] that encourage equitable load distribution among experts and enhance model training stability.
@@
Scalability and Communication Overhead. As the escalating sizes of LLMs with MoE necessitate more expansive distributed systems, the imperative for efficient communication during model training becomes increasingly critical, as elaborated in Section 5.2. The trade-off between model complexity, indicated by the number of parameters, and the communication overhead represents a significant bottleneck in distributed training processes [28]. To address these challenges, it is essential to develop and implement effective strategies that enhance the efficiency of information transfer from system aspect or streamline information exchange without compromising model performance from algorithm aspect. Innovations such as DeepSpeed [66], FasterMoE [134], and ScMoE [110] are at the forefront of minimizing communication overhead. For example, the shared expert approach [66, 69, 104, 32], advancing MoE with parameter-sharing frameworks, holds promise for reducing the volume of data transmitted between distributed systems while concurrently enhancing model performance in natural language processing tasks. Such innovations are pivotal in facilitating more scalable and efficient distributed training architectures for MoE models.
@@
Expert Specialization and Collaboration. Expert specialization refers to the concept where each expert develops non-overlapping and focused knowledge. Encouraging experts to concentrate their skills on distinct sub-tasks or domains has been shown to enhance the performance and generalization of the MoE model. The prevailing strategy involves designating a select number of experts as shared ones, with the goal of capturing commonalities in knowledge and reducing redundancy among those experts that are routed dynamically [66, 38, 104, 69]. However, fostering effective collaboration among these specialized experts is an ongoing challenge. Relying solely on a sparsely computed weighted sum of outputs from the top- experts can overlook the intricate internal relationships that exist across the entire experts. Consequently, exploring new mechanisms for enhancing both the specialization and collaboration among experts is crucial for the development of more integrated and powerful MoE models.
@@
Sparse Activation and Computational Efficiency. One of the primary benefits of MoE models lies in their capacity for sparse activations, which theoretically enhances computational efficiency. Nevertheless, implementing this efficiency in practice poses substantial challenges. This is attributed to the non-uniformity of sparse operations within hardware accelerators [217, 218]. Furthermore, optimizing the balance between activating a select top- subset of experts from an entire pool of experts entails intricate coordination. This optimization is crucial for ensuring that each expert develops a specialized niche [69]. Thus, there is a pressing need for further research into hardware optimization techniques that more adeptly accommodate sparse computations. Such advancements would not only preserve the model’s capacity but could also significantly enhance the performance and efficiency of MoE models.
@@
Generalization and Robustness. MoE models have demonstrated increased computational efficiency during pretraining phases. However, there is a notable propensity for sparse MoE architectures to overfit to specific tasks or datasets, which undermines their ability to generalize effectively [36, 37, 219, 220]. To enhance the generalization and robustness of MoE models when encountering unseen data and diverse input variations, various strategies have been explored. These include regularization techniques such as dropout [36] and token dropping [37], as well as multi-task instruction tuning [219, 220]. Looking ahead, there is potential for further advancements in this challenge. Future endeavors could explore innovative regularization methods, refined multi-task learning frameworks, or the incorporation of meta-learning concepts that bolster the MoE models’ robustness and extend their generalization capabilities across an even broader spectrum of downstream tasks.
@@
Interpretability and Transparency. The inherent complexity of MoE models, coupled with their dynamic gating of inputs to specialized experts, poses significant challenges to interpretability. This becomes particularly problematic in contexts where comprehending the rationale behind the model’s decisions is essential. Enhancing the interpretability of MoE models is therefore critical, not only to facilitate a clearer understanding of their decision-making processes but also to address underlying challenges such as load balancing [28, 36, 35] and the mitigation of knowledge redundancy [104, 69]. In light of these considerations, there is a pressing need for future studies focused on the development of methods and tools that can effectively visualize and explain the behavior of individual experts within MoE models, as well as the nature of their interactions. Such advancements would significantly improve our grasp of MoE models and bolster their ongoing development, ensuring their gating decisions are transparent and trustworthy.
@@
Optimal Expert Architecture. The design of MoE architectures, encompassing the selection of network types and the quantity of experts, significantly influences the efficacy of multi-task learning across various domains. A plethora of network architectures has been adopted as experts, including LSTM [27], CNN [99, 100], FFNs (MLPs) [28, 36, 37, 64], Attention [82, 73], and LoRA [45, 114, 47]. Among these, FFNs as experts remain the most prevalent. Despite their considerable achievements, the exploration of various hybrids of network types within experts (as the distinct features processing capabilities of different network architectures), as well as the development of innovative expert architectures, remains nascent areas of research. Furthermore, the strategic allocation of a varying number of experts across different layers of the model presents an area ripe for investigation. This is due to two primary considerations: 1) different layers of the model capture semantic information at varying levels of granularity; 2) an excessive number of experts can complicate the training process and augment computational costs, while an insufficient number of experts might lead to knowledge redundancy and diminish the specialization capabilities of the experts. To navigate these challenges, the development of automated architecture search methods specifically designed for MoE models is imperative [95]. Such approaches could systematically identify optimal configurations, balancing the trade-offs between computational efficiency and the specialization of experts.
@@
Integration with Existing Frameworks. Ensuring seamless integration of MoE models into existing large language models (LLMs) is crucial for their broad adoption. It is particularly vital to enable adaptation of LLMs to MoE architecture without necessitating training from scratch, as it can significantly reduce resource consumption. Recent studies [45, 115, 114, 47, 42, 118, 48] have demonstrated the efficacy of combining Parameter-efficient Fine-tuning (PEFT) techniques with MoE frameworks, offering a promising method for incorporating MoE into established LLMs. However, these methods may compromise model performance or complicate the existing parallel strategies of pretraining and inference efforts [186]. Advancing the development of modular and plug-and-play MoE components is essential. Additionally, optimizing these components for training and deployment across diverse computing environments and hardware platforms will expand their applicability. Such advancements are expected to enhance the versatility and efficiency of MoE models, making them more accessible for a wide range of applications and platforms.